A 17-kg piece of metal displaces 2.0 L of water when submerged. Find its density.
1
Expert's answer
2019-04-22T11:05:50-0400

The volume of the piece of metal is equal to the volume of the water displaced:


"V_{piece} = V_{water displaced} = 2.0 L \cdot \dfrac{1 m^3}{1000 L} = 2 \cdot 10^{-3} m^3."

Then, we can find the density of the piece of metal from the formula:


"\rho_{piece} = \dfrac{m_{piece}}{V_{piece}} = \dfrac{17 kg}{2 \cdot 10^{-3} m^3} = 8500 \dfrac{kg}{m^3}."

Answer:

"\rho_{piece} = 8500 \dfrac{kg}{m^3}."